# Sigma: Semantically Informative Pre-training for Skeleton-based Sign Language Understanding (Under review)

### v1

@@ 1 Introduction
Sign languages (SLs) are the primary means of communication for around 70 million people with hearing or speech impairments, spanning more than 200 SLs worldwide WHO (2025); WFD (2025). SLs remain challenging for the general public to master due to the global diversity and complex structure, which encompasses rapid and intricate hand gestures, body postures, as well as facial expressions. The ultimate goal of sign language understanding (SLU) is to comprehend SLs at the levels of words, phrases, as well as sentences anytime and anywhere for the impaired community, enabling barrier-free communication for them. Achieving this goal requires the development of models capable of interpreting these visual signals in alignment with the unique linguistic structure of SLs. SLU typically comprises three core tasks: isolated sign language recognition (ISLR), which recognises sign glosses  [^1] A sign gloss is a textual label that represents the meaning of a sign sequence using a word or phrase. Hu et al. (2021a; 2023b); Pu et al. (2024); continuous sign language recognition (CSLR), which aligns unsegmented sign sequences with sign glosses Hu et al. (2021a); Zuo & Mak (2022); Fu et al. (2025); and sign language translation (SLT), which converts sign sequences into sentences Zhou et al. (2021a; 2023); Fu et al. (2025). These tasks demand both fine-grained visual recognition and strong contextual understanding.
@@
Recently, SLU research has progressively shifted from fully supervised learning toward the development of effective pre-training paradigms, commonly referred to as sign language pre-training (SLP) Zhou et al. (2023); Hu et al. (2023b); Zhou et al. (2024); Li et al. (2025); Fu et al. (2025). These methods present a promising direction by enabling models to learn transferable representations directly from SL data, thereby significantly reducing the reliance on manual annotations, such as gloss annotations, temporal boundaries or clip-level supervision. By capturing structural and temporal regularities during the pre-training stage, models gain generalizable knowledge that accelerates convergence and enhances performance on a wide range of downstream SLU tasks. Consequently, SLP serves as a foundational step toward building unified and scalable SLU frameworks. Despite their potential, current SLP-based SLU methods continue to face significant limitations.
@@
First, the lack of semantic grounding in visual representations remains a major challenge in advancing SLU. While dense geometric features in skeletal data, such as hand trajectories, body movements, and facial expressions, provide important visual cues, they often carry limited linguistic meaning. Most existing skeleton-based SLU methods focus on capturing these low-level patterns from skeletal data, treating SL primarily as a visual signal and paying little attention to the underlying linguistic structure Hu et al. (2021a; 2023b); Zhao et al. (2024b); Pu et al. (2024). Although such models may capture low-level motion patterns, they struggle to model the relationship between these geometric features and their intended semantic roles. This disconnect weakens the ability of models to produce accurate and meaningful interpretations. Addressing this issue requires enriching visual features with semantic grounding, allowing the model to understand both the appearance and the purpose of each gesture. Doing so helps bridge the gap between visual representation and language understanding, making the model capable of supporting accurate recognition and fluent translation.
@@
Second, the imbalance between local-global feature modelling remains a persistent challenge in SLU, which inherently spans both recognition and translation tasks. Accurately distinguishing subtle variations in SL gestures requires capturing fine-grained local motion patterns, while achieving coherent understanding necessitates preserving high-level global semantics. Balancing these two levels of representation is inherently difficult but critical Liu et al. (2013). Global semantic modelling plays a key role in resolving ambiguities between visually similar SL gestures, particularly in continuous streams where the boundaries of sign glosses are unclear and context determines meaning. In such cases, local features alone are inadequate. Conversely, precise local detail extraction is equally vital, as small variations in hand gestures, body postures, or facial expression can dramatically alter meaning and grammatical structure. Even minor changes in motion intensity may shift interpretation and degrade translation quality Camgoz et al. (2018). Therefore, robust SLU demands a mechanism that jointly models both local and global features in a balanced manner.
@@
Third, inefficient cross-modal representation learning remains a critical bottleneck for advancing SLU. Compared to traditional video understanding, SLU from RGB videos is more challenging because gestures and facial expressions are more intricate or rapid than general human actions or scene changes Hu et al. (2021a; 2023b). Constructing structured, semantically aligned representations from raw visual streams is difficult, as models are easily distracted by background details or appearance variations rather than focusing on the linguistic cues that carry meaning Hu et al. (2021a; 2023b); Pu et al. (2024). This inefficiency weakens the alignment between dynamic gestures and textual semantics while imposing heavy computational and storage costs, ultimately limiting the scalability of SLU and slowing progress in SL production and generation. Skeletal data offers a promising alternative to RGB videos Hu et al. (2021a; 2023b); Pu et al. (2024). It intentionally prioritises the essential spatial-temporal dynamics of SL, which are the core semantic carriers in SLU. Modest estimation variances in skeletal data can improve generalisation across diverse real-world motion patterns, and by abstracting away visual noise such as lighting, background clutter, and appearance biases, skeletal representations provide cleaner, more relevant inputs with stronger privacy guarantees.
@@ 2.1 Semantically informative visual feature
Learning semantically informative visual features from SL sequences is crucial for understanding. This is particularly important in resolving the representation density problem, where visually similar SL gestures, differing only slightly in motion or expression, tend to cluster closely in feature space Ye et al. (2024). Incorporating linguistic and contextual cues into visual representations helps mitigate feature overlap and enables the model to learn more separable and discriminative features. This can lead to improved performance in both recognition and translation tasks, especially in cases where subtle visual differences correspond to distinct meanings. Prior works such as TSPNet Li et al. (2020b), GLE-Net Hu et al. (2021c), HST-GNN Kan et al. (2022) and SignCL Ye et al. (2024) have made progress in temporal modelling, global context extraction, and multi-perspective graph-based reasoning. However, learning and embedding semantically rich visual features in a way that generalises across tasks remains an open challenge in advancing SLU.
2.2 Sign language understanding
@@
SLU has been widely studied through task-specific methods. Prior works for ISLR have applied spatial-temporal modelling to improve accuracy Hu et al. (2021b); Li et al. (2020c); Zuo et al. (2023). Recent models for CSLR address co-articulation and gloss boundary ambiguity using CTC-based or sequence-to-sequence frameworks Min et al. (2021); Hu et al. (2023e); Jiao et al. (2023). For SLT, gloss-based approaches rely on intermediate gloss annotations Camgoz et al. (2020); Zhou et al. (2021b), while emerging gloss-free methods adopt pre-training and large language models to reduce annotation requirements and improve generalisation Zhou et al. (2023); Wong et al. (2024); Gong et al. (2024). In this study, we focus on the gloss-free SLT paradigm and aim to enhance its effectiveness by learning semantically rich visual representations aligned with textual outputs. In contrast to prior task-specific methods, we propose a unified framework capable of performing all aforementioned SLU tasks. A central design motivation lies in the differing representational needs across tasks of recognition and translation.
2.3 Sign language pre-training
@@
SLP methods employ pretext tasks to learn useful representations from SL data, improving downstream performance. Self-supervised models like SignBERT Hu et al. (2021a; 2023b) use masking and reconstruction to capture visual patterns from unlabeled videos but often lack sufficient sematic grounding. To address this, MMTLB Chen et al. (2022a) introduces multi-task training across sign-to-gloss, gloss-to-text, and sign-to-text objectives, while GFSLT-VLP Zhou et al. (2023) uses contrastive learning for sign-text alignment. More recent efforts, including MSLU Zhou et al. (2024) and C2RL Zuo & Mak (2022), incorporate keypoint reconstruction and language modelling to enhance semantic representation. Despite these advances, most approaches remain task-specific, limiting scalability. Moreover, they often struggle to balance modality-specific encoding with effective cross-modal transfer, both of which are essential for developing unified and generalisable SLU systems.
@@ 3.1 Preliminaries
We use paired skeletal data and their corresponding text(s) for both the pre-training and fine-tuning stages. The text(s) are tokenised before being fed into the text encoder. The skeletal data are 2D keypoints estimated from SL videos using RTM-Pose Jiang et al. (2023). Part-specific ST-GCNs Yan et al. (2018) are used to model both joint interdependencies and motion dynamics. Following these ST-GCNs, the raw skeletal input $S^{\text{raw}}_{p}\in\mathbb{R}^{L\times N_{p}\times C}$ is projected into a compact feature $S_{p}\in\mathbb{R}^{L\times D}$, where $L$ is the sequence length,
@@ 3.2 Sign language pre-training
We initialise Sigma with pre-trained mT5 Xue et al. (2021); Li et al. (2025) to leverage large-scale text corpora for stronger visual-linguistic alignment.
@@ 3.2.1 Sign-aware early fusion mechanism
Figure 2: SignEF promotes progressive visual-linguistic interaction with parameters $W^{(x,L)},W^{(x,N)}:x\in\{q,v,\text{out}\}$}, analogous to query, value, and output projections by Vaswani et al. (2017).
A key challenge in skeleton-based SLU is aligning geometric gesture features with textual semantics. Inspired by Vaswani et al. (2017); Li et al. (2022b), we propose SignEF, which enriches SL representations by introducing cross-modal interaction at the encoding stage, fostering more expressive and semantically aligned features. Specifically, SignEF deploys cross-attention and injects textual cues into visual encoding layers, enabling the model to perform deep and structured representation learning across modalities.
@@ 3.2.2 Hierarchical alignment learning
Balancing local and global feature modelling is essential for SLU, where recognition and translation require attention to both detailed and holistic semantics. Inspired by contrastive learning Chen et al. (2020); Radford et al. (2021); Li et al. (2022a); Hou et al. (2024), we introduce HAL as a core strategy for pre-training. HAL maximises agreement between sign-text pairs at both the global and local cluster levels. The similarity is computed as:
@@
where $f$ denotes features, $l$ indicates local, and $g$ means global. Globally, we aim to align the class-token representations $s_{cls}$ and $t_{cls}$ from the sign and text encoders. These tokens are projected into a shared embedding space using projection heads $g_{s}$ and $g_{t}$, enabling the model to capture coarse-grained semantic relationships across modalities. A cluster aggregator (see Figure 4) compresses text tokens into $k$ clusters. To focus on alignment, we compute the maximum similarity between each sign embedding and all text clusters. Locally, HAL enhances fine-grained interaction by computing local cluster-wise similarity between sign embeddings $s_{i}$ and clustered text embeddings $c_{j}$.
Algorithm 1 Cluster-wise sign-to-text similarity
1:Input: Sign tokens $\{S_{i}\in\mathbb{R}^{L\times D}\}_{i=1}^{B}$, Textual clusters $C\!\in\!\mathbb{R}^{B\times K\times D}$
@@
5:  $M\leftarrow S_{i}\,C^{\top}$ $\triangleright$ Compute cosine similarity matrix $M\in\mathbb{R}^{B\times L_{i}\times K}$
6:  $R\leftarrow\max(M,\text{dim}=3)$ $\triangleright$ Row-wise operation $R\in\mathbb{R}^{B\times L_{i}}$
7:  $w\leftarrow\mathrm{softmax}(R)$ $\triangleright$ Local-level scoring $score\in\mathbb{R}^{B}$
8:  $\text{score}\leftarrow\sum_{\text{dim}=1}(w\odot R)$
9:  $\mathbf{M}^{l}_{\mathrm{s2t}}[i]\leftarrow\text{score}$
@@
Glosses serve as simplified representations of SL segments in continuous video, and the extra supervision they provide has significantly improved SLU performance. However, they come with many limitations (see Appendix G). Our aggregator promotes hierarchical alignment by approximating gloss like groupings through local feature clustering. It groups subword-level textual tokens into semantically meaningful units. For instance, “curiosity” is split into “curios” and “ity,” and “背包” into “背” and “包” by tokenisers. Both are expressed as continuous sign sequences, while each subword could be intended to align with distinct visual segments. Our aggregator dynamically merges them into more concrete phrase-level units, with the number of clusters adaptively determined by sentence structure and bounded by sentence lengths. This enables our model to preserve compositional semantics and reduce alignment errors by preventing disjoint mappings of continuous signs.
We express the computation of the local cluster-wise similarity at Algorithm 1. For brevity, Algorithm 1 only presents the local sign-to-text similarity $\mathbf{M}^{l}_{\mathrm{s2t}}$, and the local text-to-sign similarity is computed analogously, with the positions of sign tokens and textual clusters exchanged. We experiment with several different row-wise operations and local-level scoring methods in Appendix H.1 to provide a better understanding of the design choice.
HAL is adopted to align different levels of paired features from different modalities. This dual-level strategy encourages semantically meaningful and discriminative cross-modal representations. The global and local losses are computed as follows:
@@
To balance both alignment levels, we introduce a parameter $\alpha\in[0,1]$, and define the HAL loss as:
@@
Figure 3: The overview of the cluster aggregator module. It converts sub-word token embeddings into cluster-level representations by grouping tokens, mapping them with offset indices, and aggregating hidden features for semantic alignment with visual inputs (See Appendix C for details).
Figure 4: The architecture of the SGTSGT encoder, which consists of two paths: the $STM$ path injects sign features via cross-attention for semantic alignment, and the $LM$ path preserves linguistic fluency through standard transformer layers (See Appendix D for details).

### v2

@@ 1 Introduction
Sign languages (SLs) are the primary means of communication for around 70 million people with hearing or speech impairments, spanning more than 200 SLs worldwide (WHO; WFD). SLs remain challenging for the general public to master due to the global diversity and complex structure, which encompasses rapid and intricate hand gestures, body postures, as well as facial expressions. The ultimate goal of sign language understanding (SLU) is to comprehend SLs at the levels of words, phrases, as well as sentences anytime and anywhere for the impaired community, enabling barrier-free communication for them. Achieving this goal requires the development of models capable of interpreting these visual signals in alignment with the unique linguistic structure of SLs. SLU typically comprises three core tasks: isolated sign language recognition (ISLR), which recognises sign glosses  [^1] A sign gloss is a textual label that represents the meaning of a sign sequence using a word or phrase. (hu2021signbert; hu2023signbertplus; pu2024siformer); continuous sign language recognition (CSLR), which aligns unsegmented sign sequences with sign glosses (hu2021signbert; Zuo_2022_CVPR; fu2025improving); and sign language translation (SLT), which converts sign sequences into sentences (zhou2021improving; zhou2023gloss; fu2025improving). These tasks demand both fine-grained visual recognition and strong contextual understanding.
@@
Recently, SLU research has progressively shifted from fully supervised learning toward the development of effective pre-training paradigms, commonly referred to as sign language pre-training (SLP) (zhou2023gloss; hu2023signbertplus; zhou2024scaling; Li2025sign; fu2025improving). These methods present a promising direction by enabling models to learn transferable representations directly from SL data, thereby significantly reducing the reliance on manual annotations, such as gloss annotations, temporal boundaries or clip-level supervision. By capturing structural and temporal regularities during the pre-training stage, models gain generalizable knowledge that accelerates convergence and enhances performance on a wide range of downstream SLU tasks. Consequently, SLPTserves as a foundational step toward building unified and scalable SLU frameworks. Despite their potential, current SLP-based SLU methods continue to face significant limitations.
@@
First, the lack of semantic grounding in visual representations remains a major challenge in advancing SLU. While dense geometric features in skeletal data, such as hand trajectories, body movements, and facial expressions, provide important visual cues, they often carry limited linguistic meaning. Most existing skeleton-based SLU methods focus on capturing these low-level patterns from skeletal data, treating SL primarily as a visual signal and paying little attention to the underlying linguistic structure (hu2021signbert; hu2023signbertplus; zhao2024masa; pu2024siformer). Although such models may capture low-level motion patterns, they struggle to model the relationship between these geometric features and their intended semantic roles. This disconnect weakens the ability of models to produce accurate and meaningful interpretations. Addressing this issue requires enriching visual features with semantic grounding, allowing the model to understand both the appearance and the purpose of each gesture. Doing so helps bridge the gap between visual representation and language understanding, making the model capable of supporting accurate recognition and fluent translation.
@@
Second, the imbalance between local-global feature modelling remains a persistent challenge in SLU, which inherently spans both recognition and translation tasks. Accurately distinguishing subtle variations in SL gestures requires capturing fine-grained local motion patterns, while achieving coherent understanding necessitates preserving high-level global semantics. Balancing these two levels of representation is inherently difficult but critical (liu2013global). Global semantic modelling plays a key role in resolving ambiguities between visually similar SL gestures, particularly in continuous streams where the boundaries of sign glosses are unclear and context determines meaning. In such cases, local features alone are inadequate. Conversely, precise local detail extraction is equally vital, as small variations in hand gestures, body postures, or facial expression can dramatically alter meaning and grammatical structure. Even minor changes in motion intensity may shift interpretation and degrade translation quality (camgoz2018neural). Therefore, robust SLU demands a mechanism that jointly models both local and global features in a balanced manner.
@@
Third, inefficient cross-modal representation learning remains a critical bottleneck for advancing SLU. Compared to traditional video understanding, SLU from RGB videos is more challenging because gestures and facial expressions are more intricate or rapid than general human actions or scene changes (hu2021signbert; hu2023signbertplus). Constructing structured, semantically aligned representations from raw visual streams is difficult, as models are easily distracted by background details or appearance variations rather than focusing on the linguistic cues that carry meaning (hu2021signbert; hu2023signbertplus; pu2024siformer). This inefficiency weakens the alignment between dynamic gestures and textual semantics while imposing heavy computational and storage costs, ultimately limiting the scalability of SLU and slowing progress in SL production and generation. Skeletal data offers a promising alternative to RGB videos (hu2021signbert; hu2023signbertplus; pu2024siformer). It intentionally prioritises the essential spatial-temporal dynamics of SL, which are the core semantic carriers in SLU. Modest estimation variances in skeletal data can improve generalisation across diverse real-world motion patterns, and by abstracting away visual noise such as lighting, background clutter, and appearance biases, skeletal representations provide cleaner, more relevant inputs with stronger privacy guarantees.
@@ 2 Related work
Semantically informative visual feature Learning semantically informative visual features from SL sequences is crucial for understanding. This is particularly important in resolving the representation density problem, where visually similar SL gestures, differing only slightly in motion or expression, tend to cluster closely in feature space (ye2024improving). Incorporating linguistic and contextual cues into visual representations helps mitigate feature overlap and enables the model to learn more separable and discriminative features. This can lead to improved performance in both recognition and translation tasks, especially in cases where subtle visual differences correspond to distinct meanings. Prior works such as TSPNet (li2020tspnet), GLE-Net (hu2021global), HST-GNN (kan2022sign) and SignCL (ye2024improving) have made progress in temporal modelling, global context extraction, and multi-perspective graph-based reasoning. However, learning and embedding semantically rich visual features in a way that generalises across tasks remains an open challenge in advancing SLU.
@@
Sign language understanding SLU has been widely studied through task-specific methods. Prior works for ISLR have applied spatial-temporal modelling to improve accuracy (hu2021hand; li2020transferring; zuo2023natural). Recent models for CSLR address co-articulation and gloss boundary ambiguity using CTC-based or sequence-to-sequence frameworks (min2021visual; hu2023adabrowse; jiao2023cosign). For SLT, gloss-based approaches rely on intermediate gloss annotations (camgoz2020sign; zhou2021spatial), while emerging gloss-free methods adopt pre-training and large language models to reduce annotation requirements and improve generalisation (zhou2023gloss; wong2024sign2gpt; gong2024llms). In this study, we focus on the gloss-free SLT paradigm and aim to enhance its effectiveness by learning semantically rich visual representations aligned with textual outputs. In contrast to prior task-specific methods, we propose a unified framework capable of performing all aforementioned SLU tasks. A central design motivation lies in the differing representational needs across tasks of recognition and translation.
@@
Sign language pre-training SLPTmethods employ pretext tasks to learn useful representations from SL data, improving downstream performance. Self-supervised models like SignBERT (hu2021signbert; hu2023signbertplus) use masking and reconstruction to capture visual patterns from unlabeled videos but often lack sufficient sematic grounding. To address this, MMTLB (chen2022simple) introduces multi-task training across sign-to-gloss, gloss-to-text, and sign-to-text objectives, while GFSLT-VLP (zhou2023gloss) uses contrastive learning for sign-text alignment. More recent efforts, including MSLU (zhou2024scaling) and C2RL (Zuo_2022_CVPR), incorporate keypoint reconstruction and language modelling to enhance semantic representation. Despite these advances, most approaches remain task-specific, limiting scalability. Moreover, they often struggle to balance modality-specific encoding with effective cross-modal transfer, both of which are essential for developing unified and generalisable SLU systems.
@@ 3.1 Preliminaries
We use paired skeletal data and their corresponding text(s) for both the pre-training and fine-tuning stages. The text(s) are tokenised before being fed into the text encoder. The skeletal data are 2D keypoints estimated from SL videos using RTM-Pose (jiang2023rtmpose). Part-specific ST-GCNs (yan2018spatial) are used to model both joint interdependencies and motion dynamics. Following these ST-GCNs, the raw skeletal input $S^{\text{raw}}_{p}\in\mathbb{R}^{L\times N_{p}\times C}$ is projected into a compact feature $S_{p}\in\mathbb{R}^{L\times D}$, where $L$ is the sequence length,
@@ 3.2 Sign language pre-training
We initialise Sigma with pre-trained mT5 (xue2020mT5; Li2025sign) to leverage large-scale text corpora for stronger visual-linguistic alignment.
@@ 3.2.1 Sign-aware early fusion mechanism
Figure 2: SignEF promotes progressive visual-linguistic interaction with parameters $W^{(x,L)},W^{(x,N)}:x\in\{q,v,\text{out}\}$}, analogous to query, value, and output projections by (vaswani2017attention).
A key challenge in skeleton-based SLU is aligning geometric gesture features with textual semantics. Inspired by (vaswani2017attention; li2022glip), we propose SignEF, which enriches SL representations by introducing cross-modal interaction at the encoding stage, fostering more expressive and semantically aligned features. Specifically, SignEF deploys cross-attention and injects textual cues into visual encoding layers, enabling the model to perform deep and structured representation learning across modalities.
@@ 3.2.2 Hierarchical alignment learning
Balancing local and global feature modelling is essential for SLU, where recognition and translation require attention to both detailed and holistic semantics. Inspired by contrastive learning (chen2020simple; radford2021learning; li2022blip; hou2024bagformer), we introduce SignEF as a core strategy for pre-training. SignEF maximises agreement between sign-text pairs at both the global and local cluster levels. The similarity is computed as:
@@
where $f$ denotes features, $l$ indicates local, and $g$ means global. Globally, we align the class-token representations $s_{cls}$ and $t_{cls}$ from the sign and text encoders. These class tokens are projected into a shared embedding space using projection heads $g_{s}$ and $g_{t}$, enabling the model to capture coarse-grained semantic relationships across modalities. A cluster aggregator (see Figure 4) compresses text tokens into $k$ clusters, which serve as compact semantic units to correspond to the $n$ sign tokens. To focus on cross-modality alignment, we compute the maximum similarity between each sign token $s_{i}$ and all text clusters $c_{j}$. Locally, SignEF enhances fine-grained interaction by computing local cluster-wise similarity between sign features and clustered text features.
Algorithm 1 Cluster-wise sign-to-text similarity (see Figure 9 in the Appendix I.1 for visualisation)
1:Input: Sign tokens $\{S_{b}\in\mathbb{R}^{N\times D}\}_{b=1}^{B}$, Textual clusters $C\!\in\!\mathbb{R}^{B\times K\times D}$
@@
5:  $M\leftarrow S_{b}\,C^{\top}$ $\triangleright$ Compute cosine similarity matrix $M\in\mathbb{R}^{B\times N_{b}\times K}$
6:  $R\leftarrow\max(M,\text{dim}=3)$ $\triangleright$ Row-wise operation $R\in\mathbb{R}^{B\times N_{b}}$
7:  $w\leftarrow\mathrm{softmax}(R)$
8:  $\text{score}\leftarrow\sum_{\text{dim}=2}(w\odot R)$ $\triangleright$ Local-level scoring $score\in\mathbb{R}^{B}$
9:  $\mathbf{M}^{l}_{\mathrm{s2t}}[b]\leftarrow\text{score}$
@@
Glosses serve as simplified representations of SL segments in continuous video, and the extra supervision they provide has significantly improved SLU performance. However, they come with many limitations (see Appendix H). Our aggregator promotes hierarchical alignment by approximating gloss-like groupings through local feature clustering. It groups subword-level textual tokens into semantically meaningful units. For instance, “curiosity” is split into “curios” and “ity,” and “背包” into “背” and “包” by tokenizers. Both are expressed as continuous sign sequences, while each subword could be intended to align with distinct visual segments. Our aggregator dynamically merges them into more concrete phrase-level units, with the number of clusters adaptively determined by sentence structure and bounded by sentence lengths. This enables our model to preserve compositional semantics and reduce alignment errors by preventing disjoint mappings of continuous signs.
We express the computation of the local cluster-wise similarity at Algorithm 1. For brevity, Algorithm 1 only presents the local sign-to-text similarity $\mathbf{M}^{l}_{\mathrm{s2t}}$, and the local text-to-sign similarity is computed analogously, with the positions of sign tokens and textual clusters exchanged. We experiment with several different row-wise operations and local-level scoring methods in Appendix I.1 to provide a better understanding of the design choice.
SignEF is adopted to align different levels of paired features from different modalities. This dual-level strategy encourages semantically meaningful and discriminative cross-modal representations. The global and local losses are computed as follows:
@@
To balance both alignment levels, we introduce a parameter $\alpha\in[0,1]$, and define the SignEF loss as:
@@
Figure 3: The overview of the cluster aggregator module. It converts sub-word token embeddings into cluster-level representations by grouping tokens, mapping them with offset indices, and aggregating hidden features for semantic alignment with visual inputs (See Appendix D for details).
Figure 4: The architecture of the SGT encoder, which consists of two paths: theSTM path injects sign features via cross-attention for semantic alignment, and the LM path preserves linguistic fluency through standard transformer layers (See Appendix E for details).